E-night UT 18-Nov-2019

Summary

There were between 2-4 parallel activities going on for most of the night, so it can be tricky to follow what went on from the wiki log. This summary is John's view of what happened, so there could be other activities I missed.

Summary - TCS Activities

Stephen Hooper modified the TO station (obs1) to have a purely local home directory (not NFS-mounted). This new directory was also started fresh to remove 15-years of accumulated schmutz. The last of some minor issues related to this were solved during twilight (i.e. permissions for multiple copies of IRAF config/parameter/log files). The old /home/telescope (old_home) is still available on NFS.

Petr Kubanek worked on several PCS issues and released a patch to PCS for Build 2019B related to automatic unwrapping.

Summary - AO Activities

Doug Miller started working on AO activities related to testing SOUL ESM on LFBG. After the first couple hours, he got sucked into the melee of getting AGW4 to work at RFBG. The SX SOUL full-AO loop was closed continuously for 2.7 hours. I don't think we ever got to ESM proper. We did not get an OSCO measurement in twilight.

Summary - AGW4 Checkout, Alignment, and Calibration

Petr had already checked AGW4 functionality in the daytime, so the operational startup was smooth. John did the initial alignment of the on-axis wfs pupil and the guiding hotspot. Next we encountered the problem(s) that active optics for AGW4@RFBG would NOT converge with either GCS or IDL collimating. To make matters worse, the two programs diverged in different ways.

John started adjusting zernike signs and angles in IDL by hand, but this was going absolutely nowhere. Doug joined in, and we were able to identify a super-subtle IDL bug where the wfs images from AGW4 were being analyzed with the PPOL@RDG parameters even though we specified and modified RFBG parameters. IDL still wouldn't converge, so we spent an hour running IDL zernike_collect to systematically go through Z5-Z10 applied Zernikes.

Olga assisted John with aligning the guide probe to the rotator center using eyeball collimation.

Doug ran into some issues analyzing that collected Zernike data, so John started working manually through the Zernike terms with GCS collimating. At the end of the story, John got Z4-Z11,Z22 working with GCS (don't even think of using Z12-Z21), and Doug confirmed with the Zernike data that the rotator offsets from AGW-space to primary mirror-space were correct within a few degrees.

Olga collected 20-some stars worth of AGW4/LUCI2 transform data. At the very end, John found the IDL bug that he had introduced which was stopping IDL convergence even for focus terms at RFBG.

For those who like geometrical puzzles, I think I've now understood that AGW1 and AGW2 have off-axis wfs cameras that are rotated by 180 degrees. For historical reasons, the resulting inverted pattern of Zernike gains (signs) are preserved in the GCS and IDL software. When we moved AGW4 from the left side to the right side, it seemed obvious that we should use the right-side gains. That turned out to be a wrong assumption, because the AGW4 camera is not rotated by 180 like AGW2. It turns out that the left-side gains are the ones that work for AGW4@RFBG after you correct for AGW4 using the alternate readout amplifier.

Summary - MODS

Olga Kuhn ran some MODS tests. See IT 7927 and 7929.

Summary - LUCI

Olga also ran several mask acquisitions on both LUCIs to measure the rotator zeropoints. Last I heard there was still a mystery about the LUCI1 rotator zeropoint.

Summary - PEPSIPFU Faint Object Test

We handed over to Ilya for a PEPSIPFU faint object test at 10:40 UT.

List of Tasks

Top Priority Tasks:
  • DONE New local home directory for telescope@obs1 (afternoon checkout)
  • DONEPCS patch, fix TimeToLimit overwrite (0 seconds)
  • OSCO - SOUL-SX evening twilight (Doug)
  • LUCI-2
    • DONE AGW4/LUCI-2 checkout and alignment (1-2 hours) - highest priority (John)
    • DONE AGW4/LUCI-2 rotator zeropoints (0.5 hour) - highest priority (Olga)
      • 4 masks will be used, 3 have modified LMS files to add more alignment stars, will need to manually load LMS files into RTD
      • Take average of the 4 mask alignment rotation angle
      • Visibility is 19:35-02:00 hr
        • located in /home/observer/share/Service_Observing/Eng_Nov_2019/LUCI2_AGW4_RotZP - RM507-5/4
        • located in /home/observer/Service_Observing/Eng_Nov_2019/LUCI2_AGW4_RotZP - RM507-1-3
    • DONE AGW4 transform data collection (1-2 hours) - highest priority (Olga)
    • Zeropoints (clear - any seeing) - low priority - requires min 2 airmass visits - all night
      • /home/observer/share/Service_Observing/Eng_Nov_2019/LUCI_ZP/ RM507-5/4
      • /home/observer/Service_Observing/Eng_Nov_2019/LUCI_ZP/ RM507-1-3
  • ESM checks in parallel with AGW-4 work - contingent upon building ESM reconstructors (Doug)
  • PEPSI on-sky checks - High priority - to be executed after LUCI-2/AGW-4 work is completed
Lower Priority Tasks:
  • LUCI pseudo-speckle observations - 1 hour - D. Thompson - only if no other tasks can be done AND SOUL is released for LBTO to use.
    • AO/ESM on LUCI-1, (SL on LUCI-2)
    • 3 sources availble (observe only 1)
    • Requires GEIRS with SRR mode, cube-saved mode, sub-array readouts
  • LBCs - low priority - only if no other tasks to be done and/or LBCs in the beam
    • LBC superfoc sequences - requires ~1" stable seeing and can be LBCR only (contingent on results for blue which are coming)
    • lbto@obsN:~/Startup/Startup_Sep_2019/LBC
    • need a dense star field - Possible fields are those for the copointing OBs in the "Sci-like OBs" group in the OT program file LBC_Checkout_v1.6.xml.
    • Sequence - repeat several times
      • dofpia
      • superfoc at PA=0
      • superfoc at PA=180
    • PtAutoadjust software test
    • Mixed mode w/LUCI test
  • If poor conditions (poor seeing and/or thick clouds) - backup science:
    • H-band spectral library - Bino LUCI (or LUCI-1) - rise after 9pm local time - good through morning twilight
      • /home/observer/share/Service_Observing/Eng_Nov_2019/H-band_Library on RM507-5/4
      • /home/observer/Service_Observing/Eng_Nov_2019/H-band_Library on RM507-1-3
    • Arp 187 LS - Bino LUCI (or LUCI-1) - observable 11pm-3:45am (sets ~4:45am)
      • /home/observer/share/Service_Observing/Eng_Nov_2019/LocalMergers on RM507-5/4
      • /home/observer/Service_Observing/Eng_Nov_2019/LocalMergers on RM507-1-3

Preparations

AGW4 Config Files

# 20191115 - JMH setting up AGW4 config for RFBG = LUCI_R.cfg

# start with previous AGW4 config file from 2018
cp -p LUCI_L.cfg.20180914_AGW4 LUCI_R.cfg.20191115_AGW4

# save the current AGW2 configuration with an explicit name
cp -p LUCI_R.cfg LUCI_R.cfg.20180914_AGW2


# I changed LUCI_R.cfg.20191115_AGW4 in the following ways for
#    AGW4@RFBG as I prepare to install it into LUCI_R.cfg (before sky).

1) rotAngleOffs changed from 6.0 to 0.0 to match LUCI_R.cfg for AGW2
 (is this difference residual measurement error, or a real geometry difference?)

2) changed WFS_Zscalefile to LUCI_R_RM_sub11.cfg
 (filename is sided)
 (signs of Z7,Z8,Z9,Z10 are different between LFBG/RFBG)
 (Because the AGW2 wfs camera is "upside down" this file didn't work.
  During the night we created AGW4_R_RM_sub11.cfg which is identical in terms
  of signs with LUCI_L_RM_sub11.cfg.)

3) Added the missing parameter values (actually John forgot to add this for 20191181 UT)
WFS_pupil_dir_x         1    int # which direction will the pupil shift?
WFS_pupil_dir_y         -1   int # which direction will the pupil shift?
 (LFBG uses default +1,+1)
 (not immediately clear to me why LFBG/RFBG are different. ???)

4) WFS_probeRotationFactor changed from +1 to -1 for IT 5142 solution.
(now should be -1 on either LFBG or RFBG)

5) WFS_fluxSigmaFactor changed from -1.5 to -0.5 to match what we've
been using on AGW2 (likely not an important difference)

# The following remain different for AGW4 (wherever it is).
WFS_flipX true (because of alternate readout on WFS camera installed 2018)

all the hotspot and lenslet position parameters which are AGW-dependent.

Night Log 18 Nov 2019 UT

  • OSA: J Williams
  • Support: P Kubanek, S Hooper
  • Observers: JHill, DMiller
  • ISA: O Kuhn

23:00: SHooper and PKubanek working on telescope@obs1 and TCS, with JWilliams. On mods1data, OKuhn copied uncorrupted MODS images to archive (note added to IT 7927) and ran up and checked FS alignment of LUCI1 and LUCI2. Running LUCI on rm507-5 tonight. Using RTDs that came up automatically (but I am sorely tempted to click the grey RTD buttons on the observer panel... ;))

00:00: Continued work on setting up iraf on obs1... Delaying OSCO/SX.

00:50: IRAF is sending presets now. Going to a BS star. We're skipping OSCO for the moment, at least.
  • BTW, on obs1, the /home/lbto/OSCO directory is at nfs/192.168.39.20/homes/lbto/OSCO

01:05: RFBG/AGw4: the star was in the acq camera from the start. But collimation runs away...hence the work we need to do.

AGW4@RFBG Alignment

AGW4 On-axis pupil position

01:05: JHill aligning star on wfs pinhole by adjusting guiding hotspot at BS9188 on-axis.

  • See wfsc 0002 - Approximate Center of on-axis wfs pupil at 149, 223.5. (Refined later in the evening to 146.35, 236.8 )
  • Updating GCS configuration file with these numbers (for LUCI_R.cfg WFScam_hotspot_x,y). JWilliams restarting GCS Right and resending preset.

AGW4 Guiding Hotspot

01:19 JHill relaxed maxOffsetToHotspot and maxHotspotFromProbe values to 2.54 (0.54->2.54) and 2.5 (0.5->2.5), respectively, to allow some slop in position of star wrt WFS window center. See BGTechProcedures JWilliams restarted GCS Right and resent preset.

  • Seeing is 1" as measured on LFBG and DIMM (zen-corr, but am=1.067).
  • We started with guiding hotspot at 286,242 (from use at LFBG). Shifting hotspot to -X moves star right.
  • See wfsc 0055 - Current guidecam_hotspot_x,y = 264, 256 (but see below, actually was 268, 256). (Refined later in the evening to 264, 248)

AGW4 Probe Centering

01:37: Now aligning guiding hotspot with rotator center by modifying probe coordinates - see BGTechProcedures.

  • JWilliams moving right rotator.
  • circle traced out by star is centered at 131,125 (guider image 0001)
    • x2 for binning, + (15,39) for offset in limited acq readout region => circle centered at 277, 289 in physical pixels
    • guidecam_hotspot is actually at x,y = 268, 256.
    • Difference is only 09, 33. This is pretty close. We'll collimate before repeating this measurement.
01:49 JWilliams sending new preset off-axis (x=0 y=544.3)
  • Current guidecam_hotspot_x,y = 264, 256.
  • Pupil shift direction in Y goes the way JHill thought (disc in a Friday afternoon email thread).
01:59 Preset to on-axis guide star to see if we get zernikes right. Up to this point, we have not been sending Zernikes.
  • JHill says that now the confusion begins... IDL and GCS want to send Zernikes with different signs.

02:00 DMiller testing SX AO concurrently.

02:00 JWilliams testing LUCI icon on obs1 --- in the msg lot we see some Errors:
1920.02.14 21:59:00 u | user medium    | obs1.mou | Luci ONE        | RealTimeDisplay | XMLConfig.java#importFromDOM(128)                | Using default values instead of wrong or missing entries for "de.rub.astro.lucifer.operation.userInterface.observer.realTimeDisplay.ConfigAladinControl".
1920.02.14 21:59:00 u | user medium    | obs1.mou | Luci TWO        | RealTimeDisplay | XMLConfig.java#importFromDOM(128)                | Using default values instead of wrong or missing entries for "de.rub.astro.lucifer.operation.userInterface.observer.realTimeDisplay.ConfigAladinControl".
1920.02.14 21:59:00 u | error medium   | obs1.mou | Luci ONE        | RealTimeDisplay | ConfigClient.java#importDOM(296)                 | Error while trying to import "de.rub.astro.lucifer.operation.userInterface.observer.realTimeDisplay.ConfigAladinControl"!
1920.02.14 21:59:00 u | error medium   | obs1.mou | Luci TWO        | RealTimeDisplay | ConfigClient.java#importDOM(296)                 | Error while trying to import "de.rub.astro.lucifer.operation.userInterface.observer.realTimeDisplay.ConfigAladinControl"!
but Josh reports that new images appear just fine. Maybe the message refers to a more complex use of Aladin?

02:03 Stop WFSing. Then running collimate_once multiple times.

02:14 Telescope mysteriously moved. Doug noticed a huge tip and loop opened. Happened again. Maybe it is a SOUL issue and not the telescope.

02:19 reverse signs of Z5,Z6 in IDL (agw_define.pro); then reverse Z7, Z8 also.

02:28 reverse signs of Z5,Z8,Z10 only

02:30 SX/AO on a star with R=12.4 (13.7 on SOUL, AO bin 2)
  • subwin 256x256, savemode = cube
  • FeII didn't give many counts, switched to H.
  • 1-sec for 9500 counts in the peak.
  • full AO. Doug says something strange is going on. I can see that PSF is not really round in the wings.

02:34 reverse signs of Z5,Z6,Z8,Z10

Doug temporarily fixed a wierd problem where IDL was changing to RPOL params for AGW4 in station_define.pro. That eliminated substantial confusion, but AGW4 still won't collimate correctly with IDL or GCS.

02:57 Doug and John are discussing the zernike corrections sent by the IDL script - there are sign flips or something else (offset rotation angle) needed to collimate.

Time Applied Z7 Z8
  initial 127 353
03:21 +500 Z8 165 503
  +500 Z7 -23 407

03:26 change IDL sign of Z8 for a test.

03:27 Update on-axis pupil center to a better value of 149.0, 236.8

03:32 -500 Z8 then +500 Z7 --- rotation of Z7/Z8 seems to be off by mod(45 degrees). How can that be?

Time Applied Z7 Z8
  initial 168 296
03:32 -500 Z8 -187 76
  +500 Z7 55 -159
  repeat 22 -173
  -150 Z7 +150 Z8 61 -215

zernike_collect

This data was collected in /home/lbto/ on obs2, but has been moved after the fact to /home/lbto/data/20191118_rfbg/

03:51 Running zernike_collect with collimation commented out. First run had issues.

04:16 zernike_collect with Z5_10.list (500 nm amplitude) First image is wfsc 0070. Last image is wfsc 0094.
#Labels Zern     Amp      wfsc#     scale      Avg     Factor        Measured
         0         0        70       1.0       337
         5       500     71-73       1.0      -377     -1.33      -680      -354       -97
         0         0        74       1.0       309
         0         0        74       1.0      2962
         6       500     75-77       1.0      3064      0.16      3075      2997      3121
         0         0        78       1.0      2705
         0         0        78       1.0      -479
         7       500     79-81       1.0        19     25.40       -28        57        29
         0         0        82       1.0      -386
         0         0        82       1.0      -818
         8       500     83-85       1.0     -1063     -0.47      -983     -1130     -1077
         0         0        86       1.0      -658
         0         0        86       1.0        30
         9       500     87-89       1.0      -513     -0.97      -280      -642      -618
         0         0        90       1.0      -142
         0         0        90       1.0       -57
        10       500     91-93       1.0       530      0.94       547       550       491
         0         0        94       1.0       156

Doug working on trying to analyze this.

Another wierd thing is that IDL is producing the wrong sign of focus correction (Z4) in the step between raw and rotated zernikes. John verified that GCS has the correct sign. Rotating Z4 should not change the sign. (John figured out several hours later that this was caused by a bug he had induced in agw_define.pro.)

04:45 GCS zernikes blocked. IDL> gcs_wfsc_display. GCS has the right Z6 correction, but runs away in coma.

04:51 Clear Active Optics on M2

05:02 Sending +1000 Z6 (Global Offset) produces -900 Z6 correction for GCS (with +250 parasitic Z5), so the angle isn't off by as much as 22 degrees.

05:10 Measure better AGW4 guiding hotspot of 264, 248 .

Olga's MODS tests

03:25 Started to take MODS1 and MODS2 test images. On 2nd run of swapTest.pro script, the MODS2 Blue channel filter failed to move. Clicked update on GUI and, indeed, BFILTER position cannot be read, but neither can calib unit, collimator or grating
** ERROR: BFILTER BFILTER=TIMEOUT cannot read from 192.168.139.212:8012
** Abort, Retry, or Ignore? > a
on mods2data console:
 
***ERROR: blueIMCS - More than 10 successive write errors to the Blue HEB.
     Reason: Broken pipe
     Is the Blue IMCS Comtrol Unit Offline?
     blueIMCS Aborted at 2019-11-18T03:22:06.529031

05:10 MODS2 Blue is back online. See IT 7929 for details. It looks like we lost the connection or power to the Blue IEB. I could not ping IPs 192.168.139.210 (blue IEB wago), 211 (blue ieb 1), 212 (blue ieb 2) or 213 (blue IMCS). Recovered after:
  • isisCmd --mods2 m2.ie ieb_b off/on
  • stopped and restarted all services running on mods2, i.e. ie, agw, env, tcs
  • ran mods2ColdStart

Finish AGW4 probe center alignment

05:35 Align center of rotation on hotspot
  • center of rotation at 277, 289 (measured earlier in the evening)
  • hotspot 264, 248
  • Scales: 20 pix = 1" = 0.6mm
  • Need ROT (homeoffset) and RAD (ooffset).
  • to move center to the right, decrease ROT
  • to move center up, decrease RAD
  • Need to move star left and down, by -13 pixels, -41 pixels
    • -13 pix * 0.05"/pix * 0.6 mm/" * 68 cnts/mm = -26.52 cnts. We need +26.52 cnts to move center to the left.
    • -41 pix * 0.05"/pix * 0.6 mm/" * 1000 mic/mm = -1230 mic. We need +1230 mic to move center down.
  • John editing OAC file to incorporate these deltas and issued the AGw command rdwrconfig.
  • Josh restarting GCS right and resending the preset.

LUCI2 Rotator Zeropoint

06:44 Going to check the rotator zeropoint:
  • INAF mask 1, VIPB1
06:50 -several problems. We were authorized for bino LUCI and I sent this LUCI2 script, actually even before the previous preset (an aborted attempt at Arp187) finished. No complaints from TCS. But either stars are so fuzzy or the offset did not complete correctly - images 6 and 7 are nearly on top of one another. Mask to detector alignment is great, but stars are not found, possibly because of this weird offset. 07:00 twin script for sanity check with L1.
  • twin.py is not installed on rm507-5. I copied it over from obs2:/lbt/astronomy/bin, but the #! python path in that version is incorrect for rm507-5.
  • Used OT to regenerate twinned script, binoINAF_Mask1.xml
07:08:
  • L2: 09, 10 are sky, source and 11 is slit
    • L2 alignment: dX, dY, dtheta = -1.56944, -0.64834, 0.20279 with errors RA/DEC: 0.05826, 0.06386 arcsec
  • L1 - messed up because I forgot to stop subwindowing. Did so, but need to retake "sky". source (0,0) = 37, slit = 38 (wierd vertical strip), sky w mask (-10,0) = 39, 40. Finally sky position without mask is 41 (but only 1 2.51s exp) so 42 will be it.
    • L1 alignment with 42, 37 and 38 (slit): dX, dY, dtheta = 0.81340, -0.73814, 0.47959 err RA/DEC: 0.03360, 0.00835.
  • Confirmatory images: L1 = 43 (looks like something moved during exposure). L2 = 12 (sub by 10) is not too bad.
07:30 Restarted PCS (for Petr's testing). Resent preset to Mask 1 to confirm results and figure out what is up with L1 which had a huge offset. Preset failed - needed to enter correct IE/CA. Restarting GCS right also before next preset. 07:34 Resending preset. We're going to see if GCS can collimate on right. And it did get to <=800 WFE
  • L2 acq images: 13,14,15 => dx, dy, dtheta = -2.09802, -0.59533, 0.21333 err ra/dec = 0.06179, 0.0718
  • L1 acq images: 44,45,46 => dx, dy, dtheta = 0.76608, -0.82372, 0.49948 err ra/dec = 0.02853, 0.01220
  • 4 stars (all stars used) in both cases.
  • sent offsets, taking confirmatory images: 16 (L2) and 47 (L1)
  • Note I repeated the L1 alignment later using luci1.5.VIPB1.lms. After removing one star (35), I got dx,dy,dtheta=-0.0338, -0.30271, +0.05778 and an error ra/dec = 0.08, 0.038. Then, immediately after that, using luci1.120.VIPB1.lms (local) with 4 stars I got: dx,dy,dtheta=-0.009, 0.0045, 0.00045 deg with errors ra/dec of only 0.028, 0.012. Such tiny offsets are suspicious - it is as though the SW is calculating incremental offsets from the last result. Then I did what I could to reset the RTD:
reset, stop, start aladin and reload images. Now using luci1.120.VIPB1.lms (FITS) I get earlier result. with dtheta = 0.499. 07:54 I may have mistakenly used the LMS file with only 4 stars, but when I try to select the LMS with 6, then RTD overwrites my luci1.5... selection with luci1.120... 08:00 Try a different mask, KVVDSFin luci1.4.KVVDSFin.lms ID905739. Before that PCS was restarted (Petr) and Josh restored IE/CA. 08:01 Slew to this new field. Pupil too far over on WFS image (right_wfscimage000756.fits)
  • L2 acq images will be: 16, 17, 18
  • L1 acq images will be: 47, 48, 49 --- no, error. This particlar mask does not exist in LUCI1 (this was the program with different masks for L1 and L2).
  • L2 alignment:
    • Using all 6 stars for luci1.4.KVVDSFin.lms. The first result gave low errors but no target offsets appeared in box.
    • Then I tried using the local file (luci1.4...) again and got errors that are ~1":
    • 30 0.74 0.94
    • 31 0.436 0.058
    • 32 0.096 0.488
    • 33 0.6366 -0.46
    • 34 0.155 0.290
    • 35 -2.066 -1.31
    • removing star 35 and star 33 from fit gives: dx, dy, dtheta = -2.08, -2.34, 0.20963. * dtheta is L2 from these 2 masks is +0.21 deg. send. confirmatory image is 20. * 20 minus 18 * on this field, there's a wierd stray light arc. K band.
08:28 Change ZP from 177.05 to 177.25 and asked Josh to restart PCS. 08:35 Collimation with GCS is almost as good as other AGws - at the moment, it's IDL that is having some problems. 08:39 Slewing to KVVDSFin field to check RZP
  • L2: 21,22 and 23(slit)
  • L2: dtheta = 0.4167. Wrong way...
08:45 Changed ZP to 176.85 and asked Josh to restart PCS. 08:49 Slew again to field (KVVDSFin) to check
  • L2: 24, 25 and 26 (slit)
  • L2: dx, dy, dtheta = -1.266, -1.56, 0.05 with errors RA/DEDC = 0.09 and 0.09. Using stars 30, 31, 32 and 34 (rejecting 33 and 35).
  • L2: 27 confirmatory
08:55 Try again VIPB1.... no, no time for it.

More AGW4 Alignment in parallel to Olga's Rotator Zeropoint Measurements

For wfsc image 0622, IDL has the pupil in the wrong position, but GCS OK. X=23.4 Y=122.5

07:04 Set IDL x_pupil_shift_direction to -1.0, but didn't help.

07:06 Set IDL r_pupil_shift to -0.19 (We must have been using the RPOL IDL configuration value earlier.)

IDL has the correct sign for Z8, but GCS has the better amplitude. Both programs have the wrong sign for the Z7 correction.

07:34 Create AGW$_R_RM_sub11.cfg to have custom gains/signs for AGW4@RFBG in GCS.

Note that IDL has the wrong sign for Z4 after Raw Zernikes. (John figured out several hours later that this was caused by a bug he had induced in agw_define.pro.)

Use PSF Active Gains to turn off Z9, Z10 while working on other Zernikes.

07:50 I thought GCS had the wrong X pupil shift, so I changed it from -1 to +1. (This turned out to be wrong.)

2019-11-18T07:48:56.583598+00:00 tcs1 LBT_GCSR: WFS: findSpots: r_pupil_shift = 23.579253
2019-11-18T07:48:56.583615+00:00 tcs1 LBT_GCSR: WFS: findSpots: pupil_shift_x,y = 5.376701, 22.958054
2019-11-18T07:49:26.838294+00:00 tcs1 LBT_GCSR: WFS: findSpots: r_pupil_shift = 23.579253
2019-11-18T07:49:26.838312+00:00 tcs1 LBT_GCSR: WFS: findSpots: pupil_shift_x,y = -5.376701, 22.958054

Both GCS and IDL have the sign of Z10 wrong. (both are doing flipX, or say they are)

07:56 Changed AGW$_R_RM_sub11.cfg to give GCS gain for Z10, and put PSF Active Optics Gain Z10 back to 1.0.

08:04 Undo the X-pupil shift from 07:50. Change WFS_pupil_shift_x and x_pupil_shift_direction back to -1 in both GCS and IDL. This star at x=-119, y=113 makes it clear when the shift/cutout is wrong.

2019-11-18T08:04:41.680462+00:00 tcs1 LBT_GCSR: WFS: findSpots: r_pupil_shift = 31.274983
2019-11-18T08:04:41.680479+00:00 tcs1 LBT_GCSR: WFS: findSpots: pupil_shift_x,y = 27.097818, 15.615146
2019-11-18T08:05:00.087073+00:00 tcs1 LBT_GCSR: WFS: findSpots: r_pupil_shift = 31.274983
2019-11-18T08:05:00.087090+00:00 tcs1 LBT_GCSR: WFS: findSpots: pupil_shift_x,y = -27.097818, 15.615146

08:13 Changed AGW$_R_RM_sub11.cfg to give GCS gain for Z9, and put Active Optics Gain Z9 back to 1.0.

GCS seems to be converging well now. The Zernike rotation offset might be off a few degrees, but not much more than that.

Note that AGW$_R_RM_sub11.cfg now has the same pattern of signs as LUCI_L_RM_sub11.cfg. Why? Because AGW1/AGW2 are rotated with respect to each other? (yes)

Doug has the analysis of the zernike_collect data working, but the results are hard to understand (not only because it is 1AM).
 AGw#4@RFBG (values for 20191118 UT)
            agw_params[i].wfsc.rot_offset = 0.0
            agw_params[i].wfsc.pupil_x_on_axis = 146.35 ; 20191118
            agw_params[i].wfsc.pupil_y_on_axis = 236.8  ; 20191118
            agw_params[i].wfsc.x_pupil_shift_direction = -1.0 ; 20191118
            agw_params[i].wfsc.r_pupil_shift = -0.19    ; 20191118
            agw_params[i].wfsc.xy_transpose = 0
            agw_params[i].wfsc.lenslet_rotation = 0.0   ; DLM 20180907
            agw_params[i].wfsc.lenslet_x_center = 46.   ; DLM 20180907
            agw_params[i].wfsc.lenslet_y_center = 118.  ; DLM 20180907

IDL active optics is not working for AGW4@RFBG. Signs of Z4 and Z11 are wrong. I could force them, but I haven't.

The mystery of the signs is explained by John later in the night (10:35 UT).
Doug,
        I fixed IDL for having wrong signs for Z4 and Z11 at RFBG.
It was completely my fault.  I had hardwired the wfsc.zernike_factor vectors for RFBG
in the middle of the night months ago (in agw_define.pro), and then I failed to
multiply by the right_gregorian sign vector afterwards.  I didn't collimate
with it after the fix, but I'm sure it will converge now.

Like GCS, the special signs for AGW4@RFBG seem to map the
difference between left_bent_gregorian_sign and right_bent_gregorian_sign , but those are the same in IDL ??

AGW4/LUCI2 Transform Data Collection

09:00 Take transform data on Stone_Em135_0.list
  • iif_register, /TCS
  • iif_register, RFBG
  • iif_preset, Stone_E, MODE=active, GS=0, TS=0, PA=-135, /Right
  • LUCI configuration
    • Changing to H. This star (0) has 10<Vmag<12... maybe too bright.
    • N3.75,
    • N30 Field Stop
    • Lookup table ON
  • 2.51s test image. 28, but probe was in the way.
  • 29 will be a test image with probe homed (nope, a track mode preset with gs=ts=0 still has the probe in the way)
  • 31 finally shows the star, but it's way too bright for H.
  • 32 will be a preset to star 1. It would need 4 sec and also has some companions, which can make analysis more difficult
  • 34 try star 3 and minimum exptime, 2.51. It's pretty bright, but we're running out of time and Doug OKs it after the first few images.
  • transform_collect, exp=2.51 (~/data/rfbg/Stone_Em135_0.list with edit to move star 3 to top of list)
luci2 gs #
36   star is close to sat, 56000
37   background
38 70 54k in the peak
39 9 57k, (1028,1046)
40 65 57k
41 21 53k
42 27 49k
43 85 50k
44 61 58k
45 45 58k
46 30 57k
47 36 57k
48 82 59k
49 80 58k
50 53 58k --- before this, failed for gs 77 and gs 37 and had to restart
51 4  
52 48 gs 48 is a binary
53 84 58k
54 43 59k seeing is 0.66"
55 20 57k
56 64 57k
57 89 58k
58 96 53k
59 3 finished with the image after on-axis collim
60-62   2.51 darks

10:39 UT Handover telescope to Ilya for PEPSIPFU faint object tests.

-- BarryRothberg - 12 Nov 2019
Topic revision: r22 - 27 Nov 2019, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback